Operation Summary for December 10, 2010 - December 13, 2010 |
The Integrated Luminosity for the period from 12/6/10 to 12/13/10 was 53.02 Pb-1 (inverse picobarns). NuMI reported receiving 7.1E18 protons on target during this same period. Friday Dec 10 The day shift began with Operations monitoring the Tevatron (TeV) store 8351 (158.34E30) and the Recycler stash (2.3E10), with the Antiproton Source (Pbar) stacking (43.9E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-994 (JASMIN) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported that the NuMI target helium flow came into alarm; operators notified NuMI personnel. Booster personnel configured their RF system to operate with one pre-pulse instead of two. Pbar personnel accessed their rings to change out the bulk power supply for the #3 & 4 Debuncher cooling bands. RF personnel repaired two RF stations (MIRF05 and MIRF17). Also, Electrical Support (EE) personnel brought the MI-60L power supply back into the circuit. JASMIN personnel, accompanied by safety personnel, accessed MO1 to install and then, 30 minutes later, remove film in the MTest beamline. During the evening shift, at 10:45 PM, store 8351 quenched at TeV sector E1 due to an A49 separator spark. Quench recovery began. Operators and Linac personnel addressed stability issues with the I- Source. Saturday Dec 11 The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (332.9E10), with Pbar stacking (27.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, after the E1 cryo system recovered, operators conducted a dry squeeze and then began orbit checks. Operations put the TeV into HEP shot setup at 3:30 AM and established store 8355 at 4:40 AM with an average initial luminosity (AIL) of 308E30. The I- Source output continued to vary from its nominal. During the day shift, Operations monitored the accelerator complex. During the evening shift, Operations terminated store 8355 at 4:15 PM, put the TeV into shot setup, and established store 8356 at 6 PM with an AIL of 327E30. Operations reported that an I- Source spark caused instabilities that required Linac personnel to come in. Operators brought Pelletron corrector magnet power supply fluctuations to the attention of Electron Cooling (ECool) personnel; they advised normal running for now. Operations reported that a TeV RF station (TRF6) tripped off shortly after the store went in. Sunday Dec 12 The midnight shift began with Operations monitoring store 8356 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8356 at 7:45 AM and put the TeV into shot setup. During the day shift, Operations established store 8357 at 9:05 AM with an AIL of 310E30. Operations reported that the I- Source vacuum pump failed and caused about 1.5 hours of downtime while Linac personnel came in and switched over to the H- Source. Operation reported that TRF5 tripped off and caused a momentary problem with colliding beam positions. During the evening shift, Operations terminated store 8357 at 9 PM. Operators prepared CDF and D0 for one-hour collision hall accesses. Operators tried to reset and turn on TRF5 but to no avail. Operations contacted RF personnel; they came in and replaced a bad fiber module in the RF station. Monday Dec 13 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations began TeV orbit checks at 1 AM, put the TeV into shot setup at 2:30 AM, and established store 8361 at 4:45 AM with an AIL of 324E30. Operations reported that Pelletron regulation problems delayed shot setup; operators consulted with ECool personnel and with their help reloaded corrector magnet power supply settings. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 12/13/2010 email Fermilab |
| Security, Privacy, Legal | |